What is the mass of 2.80 ×10

22

×1022 formula units of NaOH

NaOH (Molar mass = 40.0 g/mol

g/mol)


1
Expert's answer
2022-03-22T04:12:05-0400

Solution:

One mole of any substance contains 6.02×1023 formula units.

Therefore,

Moles of NaOH = (2.80×1022 formula units) × (1 mol / 6.02×1023 formula units) = 0.0465 mol

Moles of NaOH = 0.0465 mol


The molar mass of NaOH is 40.0 g/mol

Therefore,

Mass of NaOH = (0.0465 mol NaOH) × (40.0 g NaOH / 1 mol NaOH) = 1.86 g NaOH

Mass of NaOH = 1.86 g


Answer: The mass of NaOH is 1.86 grams